Mr Arnold bought 2 watches and 4 pens for $586. Mr Smith bought 1 watch and 3 pens for $361. Find the cost of the pen.
Answer by chitra(359)

Let "w" be the watch & "p" be the pen representations.

By the given data:

2w + 4p = 586 ------------------> Arnold
1w + 3p = 361 ------------------> Smith

Solving both of the equations simultaneously, we find the value of the pen & the watch.

2w + 4p = 586
2w + 6p = 722 multiplied by (2)
================ Subtracting them we get:
- 2p = -136

==> p = 68

So, by back substitution we can find the value of w, watches.

1w + 3(68) = 361
1w = 361 - 204
w = 157

thus the cost of one pen is $68 and one watch is $157

thus the solution.

regs